OpenStudy (anonymous):

Convert the angle 30°37'11" to decimal degrees and round to the nearest hundredth of a degree. please explain completely how you got to it

OpenStudy (anonymous):

step by step please <3

OpenStudy (anonymous):

IT would be be 30.62 degrees. You start by dividng 37 by 60. and 11 by 3600. Let me further explain

OpenStudy (anonymous):

there are 60 minutes in a degree and sixty seconds in a minute, so there are 3600 seconds in a degree. 30 degrees + 37 minutes + 11 seconds is 30+37/60+11/3600 degrees.

OpenStudy (anonymous):

It seems as if @ccappiello beat me the explanation.

OpenStudy (anonymous):

Degrees as it is. 1 degree is 60 minutes. So 37 minutes would be 37/60 degrees. 60 seconds is 1 minute. so 11 seconds is 11/60 minutes. 11/60 minutes is (11/60)/60 seconds

OpenStudy (anonymous):

ok so i would 1st take the 30 degree

OpenStudy (anonymous):

set it to 60 which would be

OpenStudy (anonymous):

1800

OpenStudy (anonymous):

wait im a little confused

OpenStudy (anonymous):

Use a calculator to find 37/60. Add (11/60)/60 to this result

jimthompson5910 (jim_thompson5910):

60 minutes = 1 degree 1 minute = 1/60 degrees 37 minutes = 37/60 degrees -------------------------------------- 60 seconds = 1 minute 3600 seconds = 60 minutes ...multiply both sides by 60 3600 seconds = 1 degree 1 second = 1/3600 degrees 11 seconds = 11/3600 degrees -------------------------------------- 30°37'11" = 30 degrees, 37 minutes, 11 seconds 30°37'11" = 30 degrees + 37 minutes + 11 seconds 30°37'11" = 30 degrees + 37/60 degrees + 11/3600 degrees 30°37'11" = (30 + 37/60 + 11/3600) degrees 30°37'11" = (30 + 0.61666666666667 + 0.00305555555556 ) degrees 30°37'11" = ( 30.6197222222222) degrees 30°37'11" = 30.62 degrees

OpenStudy (anonymous):

i got it im sorry i was confused i watched the video thanks :)

OpenStudy (anonymous):

i didnt understand the degrees then the min now i got :)))

jimthompson5910 (jim_thompson5910):

glad it's making sense now

OpenStudy (phi):

notice that this is almost like minutes and seconds used in time (probably not a coincidence, because the Babylonians did it this way thousands of years ago..) you must know that 30 minutes is 1/2 an hour. 30 minutes (of angle) are 1/2 a degree 15 minutes (of time) is a quarter hour (1/4 of an hour) and 15 minutes (of angle) is 1/4 of a degree (or 0.25 degrees using decimals)